OpenStudy (anonymous):

11/15=m/10...solve for m

OpenStudy (anonymous):

110/15 = 22/3

OpenStudy (anonymous):

multiply both sides of the equation by 10 and you get \[\frac{11}{15}\times10=m\]

OpenStudy (anonymous):

no i need to know what m itself is in this proportion

OpenStudy (anonymous):

which will give you 22/3 if you simplify

OpenStudy (anonymous):

oh depends what kind of question it is. Generally in math problems m is the letter given for the slope of a line.

OpenStudy (anonymous):

oh disregard the last answer. Anyways off to bed night

OpenStudy (anonymous):

@ zucchini these guys just need those answers which is their in the text book